It costs money to keep the lake at NVC clean. The cost C (in hundreds of dollars) of removing p percentage of pollution is
Answer #1


C(P)= 70p 100-P 1 if p= 100% = 100 thin C(1) = 70(1) - 70 = 10.707 hundreds of Dollary 99 150-If any